# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ Abstract.
We prove that an apparent horizon can form as a result of Einsteinian evolution in pure vacuum spacetime starting from regular initial data free of apparent horizons due to pure boundary effects. We adapt a Cauchy-double-null framework and use the boundary generalized mean curvature condition for the existence of an interior apparent horizon imposed by the author S-T Yau in [38]. In particular, we prove that the condition of [38] can be met dynamically starting from a configuration that does not verify the same through a focusing mechanism. This is the first part of a two-part sequence, and in the sequel, we will focus on explicitly constructing the Cauchy data.
1. Introduction
In this article, we consider the $3+1$ dimensional pure vacuum Einstein’s equation and investigate the issue of large data semi-global existence and dynamical existence of an apparent horizon. Consider a $3+1$ dimensional globally hyperbolic $C^{\infty}$ connected oriented Lorentzian manifold $(\mathcal{M},g)$. The vacuum Einstein equations correspond to the vanishing of the Ricci curvature of $(\mathcal{M},g)$
@@ 1. Introduction
and therefore the free gravity is described by the Weyl curvature components. The formation of a black hole is one of the central issues in general relativity and deserves deep attention from a rigorous analytic perspective. We briefly recall the following historical note
1.1. Historical background
Apparent horizons have been an object of significant interest within the classical theory of General Relativity for almost sixty years. After Schwarzschild’s discovery of his eponymous metric in 1915, it took almost twenty years before researchers came to realize the existence, within it, of a region $\mathcal{B}$ with the following surprising yet salient features: First of all, observers situated inside $\mathcal{B}$ cannot send signals to observers situated at an ideal conformal boundary at infinity, called $\mathcal{I}^{+}$. Furthermore, any observer located inside $\mathcal{B}$ lives only for finite proper time [^2] Sbierski [32] moreover showed that the termination of the observer’s proper time manifests in a particularly ferocious way, as they, in fact, get torn apart by infinitely strong tidal forces. (geodesic incompleteness). The characteristics of this region (which later came to be known as a black hole) took most of the researchers of the time aback. The consensus seemed to be that these observed phenomena have to be accidents; pathologies, only present because of the strong (spherical) symmetry inherent in the Schwarzschild solution and that, in general solutions to the Einstein equations [^3] The meaning of this phrase was not rigorous at the time, as the setup for the initial value problem in General Relativity had not yet been discovered., such phenomena would not arise. However, in the 60’s, this belief was spectacularly falsified by Roger Penrose through his celebrated incompleteness theorem [^4] This theorem, in fact, was the main reason why he was awarded the Nobel Prize in Physics back in 2020, ”for the discovery that black hole formation is a robust prediction of the general theory of relativity”.. It was Penrose [29] who introduced the notion of a trapped surface, without which one cannot state his eponymous, celebrated, incompleteness theorem :
@@ 1.1. Historical background
Closely related to the trapped surface is the notion of apparent horizon. The definition of an apparent horizon differs from that of a trapped surface by the fact that the trace $\mbox{tr}\chi$ of the null outgoing second fundamental form $\chi$ vanishes point-wise, while $\mbox{tr}\underline{\chi}$, the trace of the null incoming second fundamental form, is point-wise negative. The formal definition is as follows
@@
Given a $(3+1)$- dimensional Lorentzian manifold $(\mathcal{M},g)$, a closed spacelike 22-surface $S$ is caled apparent horizon if the fundamental forms $\chi$ and $\underline{\chi}$ have everywhere zero and negative expansions on $S$, respectively i.e.,
@@
Formally, an apparent horizon can be interpreted as the outermost boundary of a domain containing closed trapped surfaces in a Cauchy slice.
@@
The existence of an apparent horizon (and trapped surfaces contained in it) is a stable feature in the context of dynamics. Indeed, sufficiently small perturbations of Schwarzschild initial data must also contain such surfaces, by Cauchy stability. As such, incompleteness is not an accident, but rather a recurring theme in the dynamics of the Einstein equations.
At the time, the existence of an apparent horizon in a spacetime was too strong an assumption to begin with. In fact, the only way back then to guarantee its existence was to assume it at the level of initial data, but this itself can be a highly non-trivial question. The first apparent horizon formation result at the level of initial data was given in [37]. Later, [38] provided a stronger result by proving the existence of an apparent horizon at the level of the initial data set because of the boundary effect. In particular, this result did not require the inclusion of matter density (in fact, matter density could be allowed to be negative). This, however, begs the question of whether apparent horizons are dynamical objects, meaning whether they can be formed in evolution starting with data devoid of such surfaces. This problem bears high physical significance and serves as a test of reality of black holes, in the following sense. The mathematical definition of a black hole region would be without physical meaning if it did not accurately capture what physicists perceive as black holes (this is more meaningful now than ever, as scientists recently succeeded in capturing the first-ever image of a black hole). Hence, if "mathematical" black holes describe "physical" black holes, they should mathematically verify certain physical properties, one of which is dynamical formation.
The first results along this direction were obtained by Christodoulou for the Einstein equations coupled to a massless scalar field in spherical symmetry. Through a series of works [16], [14], [15], and [18], Christodoulou managed to not only prove trapped surface formation, but to understand the picture of gravitational collapse in its entirety for the given model and under the given symmetry. The breakthrough in the absence of symmetry came in [17] by the same author, and this study was, in a sense, closely related to the earlier work of S-T Yau [38] and a motivating point of this current study. In this work, Christodoulou introduced a hierarchy of small and large components in the initial data which (almost) persists under the evolution of the Einstein equations. He termed his method the short pulse method. After Christodoulou, the work [22] by Klainerman-Rodnianski reduces the size of Christodoulou’s work from about 600 to approximately 120 pages by using a slightly different hierarchy. Moreover, it reduces the number of derivatives of curvature required to prove semi-global existence from two to one. A few years later, An [3] introduces the signature for decay rates $s_{2}$ on his way to proving an extension of [22] from a finite region to a region close to past null infinity. In 2014, An and Luk [6] proved the first scale-critical trapped surface formation criterion for the vacuum equations in the absence of symmetry. While Christodoulou’s data in [17] were large in $H^{1}(\mathbb{R}^{3})$, An and Luk give data which only have to be large in $H^{\frac{3}{2}}(\mathbb{R}^{3})$, which is a scale-critical norm for the initial data. Taking advantage of the scale criticality in [6], An [2] constructs initial data that give rise not merely to trapped surfaces, but an apparent horizon, a smooth 3-dimensional hypersurface consisting of marginally outer trapped surfaces. In 2019, An [1] produces a 55-page proof of trapped surface formation for the vacuum equations, making use of the signature for decay rates and obtaining an existence result from a region close to past null infinity. In [4], An and the first author extend [1] to the case of the Einstein-Maxwell system. Later, the present authors proved the large data scale-invariant semi-global existence and trapped surface formation result in [4]. Several other studies exist in the context of Einstein-Yang-Mills [11], Einstein-Vlasov system [21, 7, 8, 9, 12], Einstein-Scalar field system, and Einstein-Spinor field system. Recently, [13], in a fundamental study, provided a short proof of the formation of a trapped surface in geodesic foliation.
@@
In this article, we focus on the study by S-T Yau [38] and prove that indeed an apparent horizon can form in a dynamical manner starting from a regular configuration. This is a new idea in the sense that we want to obtain the existence of an apparent horizon as a result of boundary effect in a Cauchy slice that is a product of evolution according to the vacuum Einstein equations 62. Let us make this idea precise. Let $\widehat{\mathcal{M}}_{t}$ be a domain with boundary $S_{t}$ in the Cauchy slice $\mathcal{M}_{t}$. On the Cauchy slice $\mathcal{M}_{t}$, the Einstein constraint equations
@@
are verified, where $h$ is the induced metric on the slice $\mathcal{M}_{t}$ and $k$ is second fundamental form of this slice. Here $\text{Scal}[h]$ denotes the scalar curvature of the metric $h$. The boundary $S_{t}$ of the domain $\widehat{\mathcal{M}}_{t}\subset\mathcal{M}_{t}$ is co-dimension $2$ in the spacetime $(\mathcal{M},g)$ and hence possesses a time-like and a spacelike second fundamental form. Let $(e_{T},e_{S})$ be the orthonormal pair spanning the tangent bundle of $S_{t}$ where $e_{T}$ is time-like and $e_{S}$ is spacelike. We define the spacelike and timelike second fundamental form of $S_{t}$ as
@@
where $\{e_{A}\}_{A=1,2}$ are the orthonormal frame tangential to $S_{t}$. If the induced metric on $S_{t}$ is denoted by $\Sigma_{AB}$, then $H=s_{AB}(\Sigma^{-1})^{AB}$ is the spacelike mean curvature of $S_{t}$ while $\kappa:=k_{AB}(\Sigma^{-1})^{AB}$ is the time-like mean curvature of $S_{t}$ also same as the restriction of the trace of the second fundamental form $k_{ij}$ of $\widehat{\mathcal{M}}_{t}$ to $S_{t}$. In addition, we also recall the notion of radius of $\widehat{\mathcal{M}}_{t}$ defined by [38]
Definition 3.
Given a region $\Omega$ and a Jordon curve $\Gamma\subset\partial\Omega$ which bounds an embedded disk in $\Omega$ and a subdomain in $\partial\Omega$, let $R_{\Gamma}$ be the supremum of $r>0$ such that $\Gamma$ does not bound a disk inside the tube of $\Gamma$ with radius $r<R_{\Gamma}$. The radius Rad of $\Omega$ is defined to be the supremum of all such $R_{\Gamma}$.
With these geometric-topological notions, [38] proves a remarkable theorem on the existence of an apparent horizon on a Cauchy slice. More precisely, the theorem is as follows
Theorem 1.2.
[38] Let $M$ be a space-like hypersurface in a spacetime. Let $g_{ij}$ be its induced metric and $k_{ij}$ be its second fundamental form. Assume that the spacelike mean curvature $H$ of $\partial M$ is strictly greater than its time-like mean curvature $\mbox{tr}_{\partial M}k=\kappa$. Let $c:=\min_{\partial M}\bigg({H-|\kappa|}\bigg)$. If $c\geq\frac{3\pi}{2\text{Rad}(M)}$, then $M$ must admit an apparent horizon in its interior.
Note that this is purely a boundary effect in the sense that if the (generalized) boundary mean curvature of a domain in the Cauchy slice is significantly higher, then an apparent horizon must exist inside. This leads to the following natural question: Can one start from a regular configuration free of any apparent horizons and from an apparent horizon in an evolutionary manner in finite time. This statement needs to be made more precise, such as exactly what it means by a regular configuration, etc. We will do this momentarily. First, compare this to the work by [17]. Working in a double null framework, [17] proves that if the incoming gravitational radiation at past null infinity is sufficiently large, then a trapped surface will form in the “interior" as a result of Einstein (pure vacuum) evolution. Since the trapped surface would need to verify $\mbox{tr}\chi<0,\mbox{tr}\underline{\chi}<0$, with enough regularity, one would naturally expect to have $\mbox{tr}\chi=0,\mbox{tr}\chi<0$ (apparent horizon) before formation of a trapped surface. In this work, we indeed adapt a different approach and a different data type that yields the formation of an apparent horizon as a result of evolution.
2. Geometric setup and semi-global development
@@
Let $(\mathcal{M},g)$ be a smooth, time-oriented, four-dimensional Lorentzian manifold. We assume that $(\mathcal{M},g)$ admits a smooth double null foliation in the following sense: there exist smooth optical functions
@@
We assume that $S_{u,\underline{u}}$ is a smooth, embedded, spacelike 22–surface diffeomorphic to $\mathbb{S}^{2}$, and we denote by $\not{\gamma}_{u,\underline{u}}$ the Riemannian metric on $S_{u,\underline{u}}$ induced by $g$. When convenient, we abbreviate $\Sigma_{u,\underline{u}}:=(S_{u,\underline{u}},\not{\gamma}_{u,\underline{u}})$.
@@
where $u_{\infty}\in\mathbb{R}$ is fixed. The hierarchy and size of the data that we impose on $H_{u_{\infty}}\cup\underline{H}_{0}$ differ in an essential way from those considered in [1, 5, 6, 11, 12]. In particular, we allow large (nonperturbative) incoming radiation along one null direction, while retaining only a degenerate smallness along the transverse direction. This constitutes the first principal new feature of the present work.
@@
$\mathcal{D}_{a,\epsilon}:=[u_{\infty},-a]\times[0,\epsilon]\times S,$
@@
$u_{\infty}\leq u\leq-a,\qquad 0\leq\underline{u}\leq\epsilon,$
and the angular variables range in $S$. Hence $\mathcal{D}_{a,\epsilon}$ is the spacetime region covered by the portion of the double null foliation determined by these bounds. A schematic representation of $\mathcal{D}_{a,\epsilon}$ is provided in Figure 1.
@@
Fix $a\gg 1$ as above and consider the slice $\mathcal{M}_{t=-a}$. We define the intersection of this slice with the semi–global development by
$\widehat{\mathcal{M}}_{-a}:=\mathcal{M}_{t=-a}\cap\mathcal{D}_{a,\epsilon}.$

### v4

# Dynamical Formation of Black Holes due to Boundary Effect in Vacuum

@@ Abstract.
We prove that a marginally outer trapped surface (MOTS) can form as a result of Einsteinian evolution in pure vacuum spacetime starting from regular initial data free of MOTSs due to pure boundary effects. We adapt a Cauchy-double-null framework and use the boundary generalized mean curvature condition for the existence of an interior MOTS imposed by the author S-T Yau in [44]. In particular, we prove that the condition of [44] can be met dynamically starting from a configuration that does not verify the same through a focusing mechanism. In fact, a very mild incoming radiation can cause a large enough generalized boundary mean curvature of an isotropically large domain so that a MOTS exists in the interior. This is fundamentally different from black hole formation by standard “gravitational collapse" and can be interpreted as the dynamical realization of a long-suspected idea in GR: MOTS can form because of “global geometry", not just quasi-local concentration of gravity/matter.
1. Introduction and the Main Theorem
In this article, we consider the $3+1$ dimensional pure vacuum Einstein’s equation and investigate the issue of large data semi-global existence and dynamical existence of an MOTS. Consider a $3+1$ dimensional globally hyperbolic $C^{\infty}$ connected oriented Lorentzian manifold $(\mathcal{M},g)$. The vacuum Einstein equations correspond to the vanishing of the Ricci curvature of $(\mathcal{M},g)$
@@ 1. Introduction and the Main Theorem
and therefore the free gravity is described by the Weyl curvature components of the spacetime Riemann curvature. In the context of these equations, the formation of a black hole is one of the central issues and deserves deep attention from a rigorous analytic perspective. We briefly recall the following historical note
1.1. Background
Marginally outer trapped surfaces (MOTSs) have played a central role in General Relativity since the mid–twentieth century. Although Schwarzschild’s solution was discovered in 1915, its global causal structure was understood only decades later, when it became clear that it contains a region $\mathcal{B}$ from which no signal can reach future null infinity $\mathcal{I}^{+}$, and in which all timelike observers encounter geodesic incompleteness in finite proper time. [^2] See Sbierski [37] for a sharp formulation showing divergence of tidal forces along incomplete timelike geodesics.
For many years, these features were widely regarded as artifacts of the high degree of symmetry of the Schwarzschild metric, rather than as generic consequences of the Einstein equations. [^3] At that stage, a rigorous formulation of the Einstein initial value problem was not yet available. This view was decisively overturned by Penrose’s incompleteness theorem in the 1960s, which showed that spacetime singularities arise under minimal and physically natural assumptions. Central to this breakthrough was Penrose’s introduction of trapped surfaces [34], a quasi-local geometric notion encoding the irreversible focusing of null geodesics. Modern developments place marginally outer trapped surfaces at the heart of this theory, providing the natural boundary between dispersive and trapped gravitational dynamics.
@@ 1.1. Background
Closely related to the trapped surface is the notion of MOTS. The definition of an MOTS differs from that of a trapped surface by the fact that the trace $\mbox{tr}\chi$ of the null outgoing second fundamental form $\chi$ vanishes point-wise, while $\mbox{tr}\underline{\chi}$, the trace of the null incoming second fundamental form, is point-wise negative. The formal definition is as follows
@@
Given a $(3+1)$- dimensional Lorentzian manifold $(\mathcal{M},g)$, a closed spacelike 22-surface $S$ is caled MOTS if the fundamental forms $\chi$ and $\underline{\chi}$ have everywhere zero and negative expansions on $S$, respectively i.e.,
@@
Formally, a MOTS can be interpreted as the outermost boundary of a domain containing closed trapped surfaces in a Cauchy slice.
@@
The existence of an MOTS (and trapped surfaces contained in it) is a stable feature in the context of dynamics. Indeed, sufficiently small perturbations of Schwarzschild initial data must also contain such surfaces, by Cauchy stability. As such, incompleteness is not an accident, but rather a recurring theme in the dynamics of the Einstein equations.
@@
In this article, we focus on the study by S-T Yau [44] and proving that indeed an MOTS can form in a dynamical manner, starting from a regular configuration. This is a completely new idea in the sense that we want to obtain the existence of a MOTS as a result of boundary effects in a Cauchy slice that is a product of evolution according to the vacuum Einstein equations. The key idea is that if a three-dimensional manifold M has a boundary with strongly positive mean curvature, the effect of this mean curvature can influence the internal geometry of $M$. Let us make this idea precise. Let $\mathcal{M}_{t}$ be a domain with boundary $S_{t}$ in the Cauchy slice $\mathcal{M}_{t}$. On the Cauchy slice $\mathcal{M}_{t}$, the Einstein constraint equations
@@
are verified, where $h$ is the induced metric on the slice $\mathcal{M}_{t}$ and $k$ is second fundamental form of this slice. Here $\text{Scal}[h]$ denotes the scalar curvature of the metric $h$. The boundary $S_{t}$ of the domain $\mathcal{M}_{t}\subset\mathcal{M}_{t}$ is co-dimension $2$ in the spacetime $(\mathcal{M},g)$ and hence possesses a time-like and a spacelike second fundamental form. Let $(e_{T},e_{S})$ be the orthonormal pair spanning the tangent bundle of $S_{t}$ where $e_{T}$ is time-like and $e_{S}$ is spacelike. We define the spacelike and timelike second fundamental form of $S_{t}$ as
@@
where $\{e_{A}\}_{A=1,2}$ are the orthonormal frame tangential to $S_{t}$. If the induced metric on $S_{t}$ is denoted by $\Sigma_{AB}$, then $H=s_{AB}(\Sigma^{-1})^{AB}$ is the spacelike mean curvature of $S_{t}$ while $\kappa:=k_{AB}(\Sigma^{-1})^{AB}$ is the time-like mean curvature of $S_{t}$ also same as the trace of the restriction of second fundamental form $k_{ij}$ of $\mathcal{M}_{t}$ to $S_{t}$. In addition, we also recall the notion of radius of $\mathcal{M}_{t}$ defined by [44]. The two entities that we are interested are $H-|\kappa|$ of $S_{t}$ and the $H$-radius or Schoen-Yau radius of a domain $\mathcal{M}_{t}$ with boundary $S_{t}$. In particular, there is a sharp threshold on $H-|\kappa|$ of $S_{t}:=\partial\mathcal{M}_{t}$ in terms of $\text{Rad}(\mathcal{M}_{t})$ that allows for an existence (and non-existence) of a MOTS inside $\widetilde{\mathcal{M}}_{1}$-see theorem 2.1 for the specifics. The vital question that arises “Can one realize this condition of [44] for the existence of MOTS in the interior of a domain in a dynamical manner?" We answer affirmatively to this question in this article. Let us state the main theorem that we prove in this article
Theorem 1.2 (Main Theorem).
There exists an open class of smooth, asymptotically flat vacuum initial data for the Einstein equations, containing no trapped surfaces and no marginally outer trapped surfaces, with the following property.
The corresponding maximal Cauchy development admits a spacelike hypersurface $\Sigma_{t_{*}}$, reached in finite proper time $t_{*}>0$, that contains a compact embedded two–sphere $S\subset\Sigma_{t_{*}}$ satisfying
$\mbox{tr}\chi(S)=0,\mbox{tr}\underline{\chi}(S)<0$
that is, $S$ is a marginally outer trapped surface.
The initial data are not assumed to be small in any global norm. The formation of $S$ occurs dynamically and prior to any breakdown of the spacetime, and is driven by a boundary–induced concentration mechanism: along a suitable interior region, the generalized mean curvature
$c:=\sup_{S}\bigg(H-|\kappa|\bigg)$
crosses the Yau threshold 2.1 while the intrinsic radius of the region remains uniformly large.
In particular, the marginally outer trapped surface is not generated by forcing the outgoing null expansion to become negative, but instead arises through a dynamic realization of the Yau boundary criterion 2.1 in vacuum general relativity.
Figure 1. Schematic of the boundary mechanism yielding a MOTS through the evolution of pure vacuum Einstein equations. The main idea is to maintain the Yau’s generalized mean curvature along u=au=-a while increasing the thickness of the interior $\mathcal{M}_{1}$ uniformly in its causal future under vacuum Einstein evolution. Roughly, the mass in the isotropically large ringed region in the Cauchy slice (top view) needs to be reasonably large. The Hawking mass of each sphere foliating this region is $\gg 1$ in our study. The remarkable fact is that the interior $\mathcal{M}_{1}$ is isotropically thick. It is natural to expect a mass-length relation for the black hole formation since, due to the scale invariance of the background (Minkowski), there is no lower bound on the mass for black hole formation. The well-known expectation is that the concentration of sufficiently large mass in a suitably defined small enough domain would lead to the formation of a black hole. The main point of this study is that a large mass can be distributed within an isotropically large domain and still form MOTS.
We begin by contrasting our result with the seminal work of Christodoulou [19]. In a double null framework, Christodoulou showed that sufficiently strong incoming gravitational radiation at past null infinity—quantified through the largeness of the shear $\hat{\chi}$ in a suitable norm—forces the formation of a trapped surface in the interior purely through vacuum Einstein evolution.
The mechanism considered in the present work is fundamentally different from the gravitational collapse via the concentration of gravitational radiation (e.g., the short pulse data technique of Christodoulou [19]). We do not assume largeness of the incoming shear in the sense of [19], nor do we seek trapped surfaces generated by driving the outgoing null expansion negative. Instead, we introduce a distinct class of initial data and a different dynamical mechanism leading to the formation of a marginally outer trapped surface. The construction exploits a boundary-induced concentration phenomenon, in which the generalized mean curvature crosses the Yau threshold while the intrinsic radius of the relevant region remains uniformly controlled.
From a technical perspective, the argument rests on three complementary components: the construction of a semi-global solution to a characteristic initial value problem (where focusing of the generalized Yau mean curvature of a topological 22-sphere is achieved), the local-in-time evolution of a Cauchy problem compatible with this characteristic development, and a quantitative control of the Schoen-Yau radius under the ensuing evolution. This combination of ideas appears to be new and does not arise in previous approaches to trapped surface or MOTS formation in vacuum general relativity. A corollary of the main theorem is stated below
Corollary 1.1 (Black hole formation without trapped-surface initial data).
There exists an open class of smooth, asymptotically flat vacuum initial data for the Einstein equations that contain no trapped surfaces and no marginally outer trapped surfaces, whose maximal Cauchy development nevertheless contains an apparent horizon.
More precisely, for such initial data, the spacetime admits a marginally outer trapped surface that forms dynamically in finite time, and hence the spacetime is causal future incomplete.
It is natural to expect a mass-length relation for the black hole formation since, due to the scale invariance of the background (Minkowski), there is no lower bound on the mass for black hole formation. This issue is deeply tied to the fact that $ADM$ mass is quite rough in terms of regularity, and hence the positive mass theorem is not stable in a smooth sense (one can have black holes with arbitrarily small masses). The well-known expectation is that the concentration of sufficiently large mass in a suitably defined small enough domain would lead to the formation of a black hole. The main point of this study is that a large mass can be distributed within an isotropically large domain and still form MOTS.
The formation of the marginally outer trapped surface occurs strictly prior to any breakdown of the spacetime and does not arise from driving the outgoing null expansion negative via concentration of curvature. Instead, the mechanism is geometric: along a dynamically evolving spacelike domain $\Omega_{t}\subset\Sigma_{t}$, the generalized boundary mean curvature $H-|\kappa|$ crosses the Schoen–Yau threshold (2.1) while the intrinsic radius $\text{Rad}(\Omega_{t})$ remains uniformly controlled.
This provides the first realization, within vacuum general relativity, of a horizon formation mechanism driven by boundary geometry rather than gravitational collapse in the sense of Christodoulou. In particular, the marginally outer trapped surface arises without requiring large incoming radiation, large null shear, or short-pulse initial data. The following three are the main points that are addressed through our study, where we take this different approach of dynamical formation of the apparent horizon
(a) It decouples horizon formation from collapse.
(b) It allows for quasi-stationary or mild dynamics to trigger a horizon.
(c) It provides a rigorous PDE realization of a long-suspected idea in GR: horizons can form because of global geometry, not just quasi-local gravity or matter energy density.
From a physical point of view, this study possesses the potential to explain an important open problem in physics: the formation of supermassive black holes (mass $\approx 10^{10}M_{\odot}$). In particular, the possibility of formation through short pulse gravitational collapse as in [19] is ruled out, and our mechanism provides a potential route for such formation.
2. Overview and strategy of the proof
@@
The first part of our argument is to establish a semi-global development of regular characteristic initial data free of trapped surfaces or MOTS. Let $(\mathcal{M},g)$ be a smooth, time-oriented, four-dimensional Lorentzian manifold. We assume that $(\mathcal{M},g)$ admits a smooth double null foliation in the following sense: there exist smooth optical functions
@@
We assume that $S_{u,\underline{u}}$ is a smooth, embedded, spacelike 22–surface diffeomorphic to $\mathbb{S}^{2}$, and we denote by $\not{\gamma}_{u,\underline{u}}$ the Riemannian metric on $S_{u,\underline{u}}$ induced by $g$. When convenient, we abbreviate $\Sigma_{u,\underline{u}}:=(S_{u,\underline{u}},\not{\gamma}_{u,\underline{u}})$. To define angular coordinates on each $S_{u,\underline{u}}$ in a smooth way, we begin by defining angular coordinates on $S_{u_{\infty},0}$. Since this is a standard 2-sphere in Minkowki space, we can use the stereographic projection coordinates $(\theta^{1},\hskip 1.42262pt\theta^{2})$ on $S_{u_{\infty},0}$. We first extend this coordinate to the whole of $\underline{H}_{0}$ by insisting that $\not{\mathcal{L}}_{\underline{L}}\theta^{A}=0$ on $\underline{H}_{0}$ for $A=1,\hskip 1.42262pt2$ and then to the whole spacetime by insisting that, for all $u$, $\not{\mathcal{L}}_{L}\theta^{A}=0$, where $L$ initially starts normal to some $S_{u,0}$. As such we have established a coordinate system $(u,\underline{u},\theta^{1},\theta^{2})$ in a neighbourhood of the initial sphere. In these coordinates, the vectors $e_{3},\hskip 1.42262pte_{4}$ become
and the metric now takes the following form:
[…]
The section that maps $p\in\mathcal{M}\mapsto\left({\left.\kern-1.2pt\theta^{1}\right|_{p}},{\left.\kern-1.2pt\theta^{2}\right|_{p}}{\left.\kern-1.2pte_{3}\right|_{p}},{\left.\kern-1.2pte_{4}\right|_{p}}\right)$ is the double null gauge that we want. We begin by decomposing curvature components and Ricci coefficients with respect to the frame $(e_{1},e_{2},e_{3},e_{4})$. Let $A,B$ take values in $\begin{Bmatrix}1,2\end{Bmatrix}$.
Let us recall the definitions of the connection coefficients in this double null framework
$\zeta_{A}:=\frac{1}{2}g(D_{A}e_{4},\hskip 1.42262pte_{3}).$
The Weyl curvature components read
@@
where $u_{\infty}\in\mathbb{R}$ is fixed. In addition, the Cauchy data is provided in the interior Cauchy slice $\mathcal{M}_{1}$ as in the diagram 2. The hierarchy and size of the data that we impose on $H_{u_{\infty}}\cup\underline{H}_{0}$ differ in an essential way from those considered in [1, 5, 6, 12, 13]. In particular, we allow large (nonperturbative) radiation along one null direction, while retaining only a degenerate smallness along the transverse direction. This constitutes the first principal new feature of the present work.
@@
$D_{a,1}:=[u_{\infty},-a]\times[0,1]\times S,~D^{{}^{\prime}}:=[u_{\infty},-a-\frac{1}{a}]\times[0,1]\times S$
@@
$u_{\infty}\leq u\leq-a,\qquad 0\leq\underline{u}\leq 1,$
and the angular variables range in $S$. Hence $D_{a,1}$ is the spacetime region covered by the portion of the double null foliation determined by these bounds. A schematic representation of $D_{a,1}$ is provided in Figure 2.
@@
Fix $a\gg 1$ as above and consider the slice $\mathcal{M}_{t=-a}$. We define the intersection of this slice with the semi-global development by
$\mathcal{M}_{2}:=\mathcal{M}_{t=-a}\cap D_{a,1}.$